OpenStudy (anonymous):

The length of a rectangle, l, is six meters more than its width, w. If the perimeter of the rectangle is 36 meters, what is the length of the rectangle, in meters? A. 12 B. 6 C. 15 D. –12

OpenStudy (johnweldon1993):

okay well l = w+6....perimeter = 36..that means 2l + 2w = 36 is l = w+6...plug that into the second equation 2(w + 6) + 2w = 36 2w + 12 + 2w = 36 4w + 12 = 36 4w = 24 w = 6 now check...if w is 6...and l = w + 6...then l must equal 12....now back into the other formula 2w + 2l = 36 2(6) + 2(12) = 36 12 + 24 = 36 36 = 36 so it checks out

OpenStudy (anonymous):

What is a possible first step in solving the following equation for x? 4x – 5 + 6x = 6 A. Divide both sides by 6. B. Combine like terms. C. Subtract 5 from both sides. D. Subtract 6 from both sides.

OpenStudy (johnweldon1993):

well...what could you do? could you add 5 to both side? yeah could you combine like terms? yeah would you divide both sides by 6? i see no reason why you would would you subtract 6 from both sides? again ...no reason to your pick

OpenStudy (anonymous):

a

OpenStudy (anonymous):

no b

OpenStudy (anonymous):

like terms

OpenStudy (anonymous):

That's right.

OpenStudy (anonymous):

what

OpenStudy (anonymous):

B

OpenStudy (anonymous):

k

OpenStudy (anonymous):

At her job, Mary works 4 hours on Monday, 3 hours on Wednesday, and 6 hours on Saturday. Her boss adds a $10 bonus to her paycheck, which then totals $120.50. What is Mary's hourly pay? A. $8.50 B. $9.27 C. $5.24 D. $8.05

OpenStudy (anonymous):

C

OpenStudy (anonymous):

k

OpenStudy (anonymous):

In solving an equation for x, you distribute a –4 on one side of the equation. What needs to be done to the other side of the equation? A. multiply by 4 B. add –4 C. subtract 4 D. nothing

OpenStudy (anonymous):

D

OpenStudy (anonymous):

At his job, Rob works 6 hours on Monday, 7 hours on Tuesday, and 2 hours on Friday. He receives a $15 bonus, which causes his weekly paycheck to total $153.75. What is Rob's hourly pay? A. $8.00 B. $9.25 C. $10.25 D. $10.75

OpenStudy (anonymous):

Give me a medal for helping you.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

At his job, Rob works 6 hours on Monday, 7 hours on Tuesday, and 2 hours on Friday. He receives a $15 bonus, which causes his weekly paycheck to total $153.75. What is Rob's hourly pay? A. $8.00 B. $9.25 C. $10.25 D. $10.75

OpenStudy (anonymous):

A